OpenStudy (anonymous):

Solve -x-4(3x-4)=-62 show work

OpenStudy (anonymous):

\[-x-4(3x-4)=-62\] multiplying both sides of the equation by -1 (to make it friendlier), we have; \[x+4(3x- 4)=62\] expanding the bracketed terms; \[x+12x-16=62\] grouping like terms; \[13x=16+62\] reducing further; \[13x = 78\] making x the subject of the formula; \[x = \frac{ 78}{ 13 }\] and finally, \[x = 6\] You can check by substituting 6 for x in the left hand sid of the equation an solving to see if it equals the left hand side